OpenStudy (anonymous):

how do you graph the ellipse 2x^2+y^2=8. And how do i find the foci?

OpenStudy (anonymous):

Start by putting it into standard form. \[\frac{(x-h)^2}{a^2} + \frac{(y-k)^2}{b^2} = 1\]

OpenStudy (anonymous):

(h,k) is the point of the center. a is the length of the elipse in the x direction b is the length of the elipse in the y direction The distance from the center to the focus is \(\sqrt{a^2-b^2}\) when a is the major axis, or \(\sqrt{b^2-a^2}\) if b is the major axis.